OpenStudy (anonymous):

suppose you have $3000 to invest. which investment yields the greater return over 10 years: 6.5% compounded semiannually or 6% compounded monthly? how much more is yielded by the better investment?

OpenStudy (anonymous):

Well 10*.065 because you move the decimal two places to the left equals .65 add that to your 3000. then do your 10*.06 to get .6 then add it to your 300 i think that is how you would do this type of problem. Does it have any answers to it that you have to choose from?

OpenStudy (anonymous):

unless you would times your investment 3000 by .065 & then .06. 3000+195(10)=4950 and the second one 3000+180(10)=4800.

OpenStudy (anonymous):

\[Use the equation A=P(1+r/n)^{t*n} Where P is the amount invested r=interest rate, n is the amount of \times the interest is done a year(quarterly, semianually), t is the time invested for\]

OpenStudy (anonymous):

Use that equation P= amount invested, r=interest rate, n=amount of times the interest is compounded in(quarterly or semiannually), t is amount of time money is invested.

OpenStudy (ranga):

Compound Interest formula:\[\Large A = P(1 + \frac{ r }{ n })^{nt}\] P = Principal A = Amount at maturity (or the balance on the account) r = annual rate of interest in decimal n = compounding period t = number of years For both cases, P = 3000, t = 10 For 6.5% compounded semiannually, r = 0.065, n = 2 (semi-annual means twice a year) Find A. For 6% compounded monthly, r = 0.06, n = 12 Find A Compare and see which one gives higher A.

OpenStudy (ranga):

@shannonb0284: To keep text from bunching together when using the Equations editor try: \text { Whatever you want to stay without the words bunching up } : \[\text { Whatever you want to stay without the words bunching up } \]